40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 27 17:14:08 2015, ericq, Update, LSC, X Aux Laser crystal temperature changed 
    Reply  Fri May 29 02:05:08 2015, ericq, Update, LSC, End Laser temperatures set 
Message ID: 11335     Entry time: Fri May 29 02:05:08 2015     In reply to: 11328
Author: ericq 
Type: Update 
Category: LSC 
Subject: End Laser temperatures set 

Both green beatnotes have been found with nominal amplitudes. (X: -30dBm Y: -20dBm), at temperatures which don't seem to be prone to mode hopping. 

X = 42.64, Y ~40.15

Both arms can lock on ALS, but as Koji mentioned in ELOG 11334, the ALS noise seems anomalously high. 


Details

The temperatures I posted in the previous log ended up not being so useful. To find the right end laser temperatures, I looked at the IR beatnotes on the control room analyzer out to 1GHz, and swept around the SLOW_SERVO2_OFFSET channels to change the laser temperature. During this time, the end green shutters were closed, the PSL shutter was closed, the FSS slow servo was off, and the FSS_SLOWDC was set to 0.0. (The green PSL shutter had to be open, because the IR beat fiber is coupled after it.)

For each arm, I found three temperatures where an IR beat could be observed; as Koji mentioned on Wednesday, we should use the middle mode. For each of the arms, scanning around from the middle to move the beat by +-1GHz did not cause a mode hop - the beat stayed visible on the scope. Once I found a real IR beat for the X arm, I took at look at the RF output of the X BBPD, and found my alignment from the other night was actually pretty good; I made a minor touch up to maximize the green beat. 

For the AUX X innolight, I was able to find the actual temperature of the laser crystal when at the correct point, remove the digital offset, and return to the same temperature with the controller front panel dial. This temperature is 42.64 degrees Celcius. There is no diode temperature control on the unit, as far as I could tell, but the maximum green transmitted power through the X arm is about the same as it's ever been, around GTRX=0.5. 

My motivation for doing this was that I always have problems remembering the historical typical values for the digital temperature offset. It seems much cleaner to me to set things up such that a beat is visible "at the origin" (i.e. FSS_SLOWDC=0 and SLOW_SERVO2_OFFSET=0) I suppose that this will also depend on what mode of the PMC we're locked to. During my time working today, its been locked near the middle of its range, currently sitting at 145V. 

However, for the AUX Y lightwave, I was a little perplexed to find that moving the the laser crystal setpoint around does not apear to change the real laser temperature at all. Thus I could not offload the digital offset in the same way. Aditionally, the lightwave controller does not have the same temperature measuring accuracy as the innolight, even with the back panel voltage readout that is hooked up to the multimeter that lives under the optical table. The best I can tell is around 40.1-40.2 degrees C. Y_SLOW_SERVO2_OFFSET of -10690 counts gives a beat <100MHz at FSS_SLOWDC=0. This is actually very close to the previous operating point, so the same mode seems to still work. 

The arms now easily lock on ALS, albeit with higher noise. With arms locked on ALS, POX and POY show >1kHz rms noise. frown 


I gave the PRFPMI locking script a few tries, but it's having problems keeping the PRMI locked. The ALS is a few times noisier than usual, and I haven't revisited the validity of the PRC angular feedforward, so I'm not so surprised. 

ELOG V3.1.3-